. Simple random sampling uses a sample of size n from a population of size N to obtain data that can be used to make inferences about the characteristics of a population. Suppose that, from a population of 50 bank accounts, we want to take a random sample of four accounts in order to learn about the population. How many different random samples of four accounts are possible


1
Expert's answer
2021-05-19T12:41:51-0400
"C_{50}^4=\frac {50!}{4!\times (50-4)!}=230300"